### v1

@@ 1 Introduction
for some approximation factor $\Gamma$ [^1] If $\bm{A}$ is itself in $\mathcal{S}$, then this problem is referred to as matrix recovery. Approximation guarantees for other norms are also of interest. Moreover, we wish to do so given only implicit access to $\bm{A}$ through black-box matrix-vector product queries $\bm{x}\mapsto\bm{A}\bm{x}$ and $\bm{x}\mapsto\bm{A}^{\mathsf{T}}\bm{x}$, both of which will henceforth be referred to as “matvec queries to $\bm{A}$.” Perhaps the most well-studied version of this problem is when $\mathcal{S}$ is the class of low-rank matrices, for which methods like (block) Krylov iteration and the randomized SVD can be used [36, 6, 7, 63]. Recent work has also considered approximation algorithms for classes such as diagonal matrices [12, 64, 10, 29], sparse matrices [27, 26, 25, 69, 28, 58, 3], and beyond [68, 43, 62, 47, 37].
@@ 1.1 Hierarchical low-rank Matrices
In this paper, we specifically study algorithms for solving Eq. 1 where $\mathcal{S}$ is the widely used class of hierarchically semi-separable (HSS) matrices (see Definition 2.3). Like the broader class of hierarchical off-diagonal low-rank (HODLR) matrices, HSS matrices are recursively partitioned into low-rank off-diagonal blocks and on-diagonal blocks. However, the column and row spaces of the low-rank blocks of HSS matrices at different levels of the hierarchy satisfy an additional nestedness property. This results in a format that is even more compact than that of HODLR matrices. At the same time, the added structure makes algorithms for HSS approximation substantially harder to design and analyze. Indeed, while there has been significant work on algorithms for obtaining HSS approximations (see Subsection 1.3 for details), even when the entries of $\bm{A}$ can be accessed explicitly, we are not aware of any polynomial-time algorithms that achieve a strong multiplicative error bound as in Eq. 1. In the setting where $\bm{A}$ is only accessible through matrix-vector products, we are only aware of one existing algorithm by Levitt and Martinsson [41]. However, while the algorithm works well in practice, as noted in [41, Remark 4.3] it is not well-understood theoretically.
@@ 1.3 Past work
Most closely related to the present work is [23], which analyzes a matvec query algorithm for HODLR matrix approximation. The algorithm from [23] can be viewed as a robust version of widely used “peeling” methods for recovering matrices with exact HODLR structure from matrix-vector products [45, 51, 42]. It obtains a rank-$k$ HODLR approximation $\bm{B}$ that is within a $(1+\varepsilon)$ factor of optimal using $O(k\log^{4}(N)/\varepsilon^{3})$ matrix-vector products. Since the class of HSS matrices is a subset of HOLDR matrices, a simple triangle inequality argument implies that the best HSS rank-$k$ approximation to $\bm{B}$ is within a $(2+2\epsilon)$ factor of the optimal HSS approximation. In principal, this yields an approach to solving Eq. 1 but (ii) such an approach has worse matvec complexity than our method, which uses $O(k\log(N/k))$ queries and (ii) it is not computationally efficient, since it is not clear how to compute the best HSS approximation to the (explicit) HODLR matrix $\bm{B}$.
Also related to our work is research on algorithms for HSS approximation is the setting where we have explicit access to $\bm{A}$’s entries [13, 55, 40, 70]. Existing algorithms approximate low-rank submatrices up to a given fixed tolerance (e.g., using the truncated SVD). As such, they provide additive error bounds with respect to this tolerance. For instance, [40, Theorem 4.7] guarantees spectral norm error $O(\sqrt{N}\epsilon)$, where $\epsilon$ is the truncation tolerance for each low-rank block. As far as we are aware, such bounds do not translate to relative error bounds like those sought in Eq. 1, which compare to the best possible HSS approximation of a given rank.
@@ 2 Background and Notation
We denote the best rank-$k$ approximation to a matrix $\bm{M}$ in any unitarily invariant norm as $[\![\bm{M}]\!]_{k}$. This approximation can be obtained by the truncated SVD as described by the Eckart–Young Theorem. We write the transpose of $\bm{M}$ as $\bm{M}^{\mathsf{T}}$ and the Moore-Penrose pseudoinverse as $\bm{M}^{\dagger}$. The Frobenius and spectral norms are denoted by $\|\bm{M}\|_{\mathsf{F}}$ and $\|\bm{M}\|_{2}$, respectively. Throughout, $\operatorname{\mathbb{E}}[\,\cdot\,]$ denotes expectation and $\text{Gaussian}(m,n)$ denotes an $m\times n$ matrix whose entries are independent standard normal random variables. We denote by $\operatorname{blockdiag}(\bm{B}_{1},\ldots,\bm{B}_{d})$ the block-diagonal matrix with (possibly rectangular) diagonal blocks $\bm{B}_{1},\ldots,\bm{B}_{d}$.
@@ 2.2 Rank-structured matrices
Following [41], we define HSS matrices using a so-called telescoping factorization. [^2] For a discussion on the equivalence between the telescoping factorization and the perhaps more common “data-sparse” definition of HSS matrices, we refer the reader to [21, Section 3].
@@
Consider a matrix $\bm{B}\in\mathbb{R}^{2^{L+1}k\times 2^{L+1}k}$. We say that $\bm{B}\in\operatorname{HSS}(L,k)$ if it satisfies
@@
$\bm{U}^{(\ell)}$ $=\operatorname{blockdiag}\left(\bm{U}_{1}^{(\ell)},\ldots,\bm{U}_{2^{\ell}}^{(\ell)}\right)\quad\text{where }\bm{U}_{i}^{(\ell)}\in\mathbb{R}^{2k\times k}\text{ has orthonormal columns,}$,
$\bm{V}^{(\ell)}$ $=\operatorname{blockdiag}\left(\bm{V}_{1}^{(\ell)},\ldots,\bm{V}_{2^{\ell}}^{(\ell)}\right)\quad\text{where }\bm{V}_{i}^{(\ell)}\in\mathbb{R}^{2k\times k}\text{ has orthonormal columns,}$,
@@
We note that the telescoping factorization in Definition 2.3 encodes the recursive nature of an HSS matrix. That is, if $\bm{B}=\bm{B}^{(L+1)}$ is an $N\times N$ matrix in $\operatorname{HSS}(L,k)$, then $\bm{B}^{(L)}$ is an $N/2\times N/2$ matrix in $\operatorname{HSS}(L-1,k)$, and so on. One level ($\ell=2$) of the telescoping decomposition is illustrated below. The telescoping factorization also implies that for any $i\neq j$, the $(i,j)$ block of $\bm{B}^{(L+1)}$ is rank-$k$, with range and co-range respectively contained in the ranges of $\bm{U}^{(\ell)}_{i}$ and $\bm{V}^{(\ell)}_{j}$. This immediately implies the HSS block-rows and block-columns are rank-$k$.
=+$\bm{B}^{(\ell+1)}$=$\bm{U}^{(\ell)}$$\bm{B}^{(\ell)}$$(\bm{V}^{(\ell)})^{\mathsf{T}}$+$\bm{D}^{(\ell)}$
@@
Definition 2.3 requires that $\bm{B}$ has dimension $N=2^{L+1}k$, so $L=\log_{2}(N/k)-1.$ This is to simplify notation, and our analysis easily extends to arbitrary $N$. We discuss this and the relation to other hierarchical classes in Appendix B.
@@
$\bm{U}$ $=\operatorname{blockdiag}\left(\bm{U}_{1},\ldots,\bm{U}_{2^{L}}\right)\quad\text{where }\bm{U}_{i}\in\mathbb{R}^{2k\times k}\text{ has orthonormal columns,}$,
[…]
@@ 3 Near-optimal greedy HSS approximation
3:   for $i=1,2,\ldots,2^{\ell}$ do
4:      Partition $\bm{A}^{(\ell+1)}=\{\bm{A}^{(\ell+1)}_{i,j}\}_{i,j}^{2^{\ell}}$ into blocks of size $\bm{A}_{i,j}^{(\ell+1)}\in\mathbb{R}^{2k\times 2k}$.
@@
$\hskip 7.20003pt\leq\Gamma_{\textup{d}}\cdot\Big{[}\|r_{i}(\bm{A}^{(\ell+1)})-\bm{U}_{i}^{(\ell)}(\bm{U}_{i}^{(\ell)})^{\mathsf{T}}r_{i}(\bm{A}^{(\ell+1)})\|_{\mathsf{F}}^{2}+\|c_{i}(\bm{A}^{(\ell+1)})-c_{i}(\bm{A}^{(\ell+1)})\bm{V}_{i}^{(\ell)}(\bm{V}_{i}^{(\ell)})^{\mathsf{T}}\|_{\mathsf{F}}^{2}\Big{]}.$
Then, [^3] We are not aware of any results that guarantee the existence of an optimal HSS approximation. We thus use the infimum, instead of minimum, in (3).
(3) $\mathbb{E}\Bigl{[}\|\bm{A}-\bm{B}\|_{\mathsf{F}}^{2}\Big{]}\leq\left(\Gamma_{\textup{r}}+\Gamma_{\textup{c}}\right)(1+\Gamma_{\textup{d}})L\cdot\inf\limits_{\bm{C}\in\operatorname{HSS}(L,k)}\|\bm{A}-\bm{C}\|_{\mathsf{F}}^{2}.$
As discussed Algorithm 1 can be implemented in many different ways [41, 49, 50, 72, 71]. We highlight some important examples below:
Explicit matrices: If $\bm{A}$ is dense and stored in memory, the simplest way to implement Algorithm 1 is to use the truncated SVD: let $\bm{U}_{i}^{(\ell)}$ and $\bm{V}_{i}^{(\ell)}$ be the top $k$ left and right singular vectors of $r_{i}(\bm{A}^{(\ell+1)})$ and $c_{i}(\bm{A}^{(\ell+1)})$, respectively, and let $\bm{D}_{i,i}^{(\ell)}=\bm{A}_{i,i}^{(\ell)}$. This implementation requires $O(N^{2}k)$ operations. Since at each level we compute the optimal subspaces of each block-row and block-column, in the context of Theorem 3.1, this yields $\Gamma_{\textup{r}}=\Gamma_{\textup{c}}=1$ and $\Gamma_{\textup{d}}=0$. Consequently, for this implementation of Algorithm 1, the output satisfies $\|\bm{A}-\bm{B}\|_{\mathsf{F}}^{2}\leq 2L\inf\limits_{\bm{C}\in\operatorname{HSS}(L,k)}\|\bm{A}-\bm{C}\|_{\mathsf{F}}^{2}.$
Implicit matrices: When $\bm{A}$ can only be accessed using matvecs, we leverage the “block-nullification” idea of [41] to implement Algorithm 1. This approach is described in Section 4, where we prove in Theorem 4.1 that the output of this algorithm satisfies $\mathbb{E}\|\bm{A}-\bm{B}\|_{\mathsf{F}}^{2}\leq O(L)\inf\limits_{\bm{C}\in\operatorname{HSS}(L,k)}\|\bm{A}-\bm{C}\|_{\mathsf{F}}^{2},$ yielding our main result.

### v2

@@ 1 Introduction
for some approximation factor $\Gamma$. [^1] If $\bm{A}$ is itself in $\mathcal{S}$, then this problem is referred to as matrix recovery. Approximation guarantees for other norms are also of interest. Moreover, we wish to do so given only implicit access to $\bm{A}$ through black-box matrix-vector product queries $\bm{x}\mapsto\bm{A}\bm{x}$ and $\bm{x}\mapsto\bm{A}^{\mathsf{T}}\bm{x}$, both of which will henceforth be referred to as “matvec queries to $\bm{A}$.” Perhaps the most well-studied version of this problem is when $\mathcal{S}$ is the class of low-rank matrices, for which methods like (block) Krylov iteration and the randomized SVD can be used [38, 6, 7, 65]. Recent work has also considered approximation algorithms for classes such as diagonal matrices [12, 67, 10, 31], sparse matrices [29, 28, 27, 72, 30, 60, 3], and beyond [71, 45, 64, 49, 39].
@@ 1.1 Hierarchical low-rank matrices
In this paper, we specifically study algorithms for solving Eq. 1 where $\mathcal{S}$ is the widely used class of hierarchically semi-separable (HSS) matrices (see Definition 2.3). Like the broader class of hierarchical off-diagonal low-rank (HODLR) matrices, HSS matrices are recursively partitioned into low-rank off-diagonal blocks and on-diagonal blocks. However, the column and row spaces of the low-rank blocks of HSS matrices at different levels of the hierarchy satisfy an additional nestedness property. This results in a format that is even more compact than that of HODLR matrices. At the same time, the added structure makes algorithms for HSS approximation substantially harder to design and analyze. Indeed, while there has been significant work on algorithms for obtaining HSS approximations (see Subsection 1.3 for details), even when the entries of $\bm{A}$ can be accessed explicitly, we are not aware of any polynomial-time algorithms that achieve a strong multiplicative error bound as in Eq. 1. In the setting where $\bm{A}$ is only accessible through matrix-vector products, we are only aware of one existing algorithm by Levitt and Martinsson [43]. However, while the algorithm works well in practice, as noted in [43, Remark 4.3] it is not well-understood theoretically.
@@ 1.3 Past work
Most closely related to the present work is [25], which analyzes a matvec query algorithm for HODLR matrix approximation. The algorithm from [25] can be viewed as a robust version of widely used “peeling” methods for recovering matrices with exact HODLR structure from matrix-vector products [47, 53, 44]. It obtains a rank-$k$ HODLR approximation $\bm{B}$ that is within a $(1+\varepsilon)$ factor of optimal using $O(k\log^{4}(N)/\varepsilon^{3})$ matrix-vector products. Since the class of HSS matrices is a subset of HOLDR matrices, a simple triangle inequality argument implies that the best HSS rank-$k$ approximation to $\bm{B}$ is within a $(2+2\epsilon)$ factor of the optimal HSS approximation. In principal, this yields an approach to solving Eq. 1 but (i) such an approach has worse matvec complexity than our method, which uses $O(k\log(N/k))$ queries and (ii) it is not computationally efficient, since it is not clear how to compute the best HSS approximation to the (explicit) HODLR matrix $\bm{B}$.
Also related to our work is research on algorithms for HSS approximation is the setting where we have explicit access to $\bm{A}$’s entries [15, 57, 42, 73]. Existing algorithms approximate low-rank submatrices up to a given fixed tolerance (e.g., using the truncated SVD). As such, they provide additive error bounds with respect to this tolerance. For instance, [42, Theorem 4.7] guarantees spectral norm error $O(\sqrt{N}\epsilon)$, where $\epsilon$ is the truncation tolerance for each low-rank block. As far as we are aware, such bounds do not translate to relative error bounds like those sought in Eq. 1, which compare to the best possible HSS approximation of a given rank.
@@ 2 Background and Notation
We denote the best rank-$k$ approximation to a matrix $\bm{M}$ in any unitarily invariant norm as $\bm{M}]\!]_{k}$. This approximation can be obtained by the truncated SVD as described by the Eckart–Young Theorem. We write the transpose of $\bm{M}$ as $\bm{M}^{\mathsf{T}}$ and the Moore-Penrose pseudoinverse as $\bm{M}^{\dagger}$. The Frobenius and spectral norms are denoted by $\|\bm{M}\|_{\mathsf{F}}$ and $\|\bm{M}\|_{2}$, respectively. Throughout, $\operatorname{\mathbb{E}}[\,\cdot\,]$ denotes expectation and $\text{Gaussian}(m,n)$ denotes an $m\times n$ matrix whose entries are independent standard normal random variables. We denote by $\operatorname{blockdiag}(\bm{B}_{1},\ldots,\bm{B}_{d})$ the block-diagonal matrix with (possibly rectangular) diagonal blocks $\bm{B}_{1},\ldots,\bm{B}_{d}$.
@@ 2.2 Rank-structured matrices
Following [43], we define HSS matrices using a so-called telescoping factorization. [^2] For a discussion on the equivalence between the telescoping factorization and the perhaps more common “data-sparse” definition of HSS matrices, we refer the reader to [23, Section 3].
@@
Consider a matrix $\bm{B}\in\mathbb{R}^{2^{L+1}k\times 2^{L+1}k}$. We say that $\bm{B}\in\operatorname{HSS}(L,k)$ if there exist block-diagonal matrices
$\bm{U}^{(\ell)},\bm{V}^{(\ell)}\text{ for }\ell=1,\ldots,L,\quad\bm{D}^{(\ell)}\text{ for }\ell=0,\ldots,L$
such that
@@
$\bm{U}^{(\ell)}$ $=\operatorname{blockdiag}\left(\bm{U}_{1}^{(\ell)},\ldots,\bm{U}_{2^{\ell}}^{(\ell)}\right)\quad\text{where }\bm{U}_{i}^{(\ell)}\in\mathbb{R}^{2k\times k}\text{ has orthonormal columns,}$
$\bm{V}^{(\ell)}$ $=\operatorname{blockdiag}\left(\bm{V}_{1}^{(\ell)},\ldots,\bm{V}_{2^{\ell}}^{(\ell)}\right)\quad\text{where }\bm{V}_{i}^{(\ell)}\in\mathbb{R}^{2k\times k}\text{ has orthonormal columns,}$
@@
We note that the telescoping factorization in Definition 2.3 encodes the recursive nature of an HSS matrix. That is, if $\bm{B}=\bm{B}^{(L+1)}$ is an $N\times N$ matrix in $\operatorname{HSS}(L,k)$, then $\bm{B}^{(L)}$ is an $N/2\times N/2$ matrix in $\operatorname{HSS}(L-1,k)$, and so on. One level ($\ell=2$) of the telescoping decomposition is illustrated below. The telescoping factorization also implies that for any $i\neq j$, the $(i,j)$ block of $\bm{B}^{(\ell+1)}$ has rank at most $k$, with range spanned by the columns of $\bm{U}^{(\ell)}_{i}$ and co-range spanned by the columns of $\bm{V}^{(\ell)}_{j}$, for all $\ell=1,\ldots,L$. This immediately implies the HSS block-rows and block-columns are rank-$k$.
$\bm{B}^{(\ell+1)}$$\bm{U}^{(\ell)}$$\bm{B}^{(\ell)}$$(\bm{V}^{(\ell)})^{\mathsf{T}}$$\bm{D}^{(\ell)}$
@@
Definition 2.3 requires that $\bm{B}$ has dimension $N=2^{L+1}k$, so $L=\log_{2}(N/k)-1$. This is to simplify notation. More generally, HSS matrices are defined in terms of a binary partition tree, where each node in the tree corresponds to an index set. In Definition 2.3, we assume that the binary tree is a perfect binary partition tree with leaf nodes containing $2k$ indices. This ensures that all basis matrices $\bm{U}_{i}^{(\ell)}$ and $\bm{V}_{i}^{(\ell)}$ have dimensions $2k\times k$, and the diagonal blocks $\bm{D}_{i}^{(\ell)}$ have dimensions $2k\times 2k$, independently of $i$ and $\ell$, which simplifies notation. However, our analysis easily extends to arbitrary $N$ and diagonal block size, corresponding to arbitrary partition trees. We discuss this and the relation to other hierarchical classes in Appendix B.
@@
$\bm{U}$ $=\operatorname{blockdiag}\left(\bm{U}_{1},\ldots,\bm{U}_{2^{L}}\right)\quad\text{where }\bm{U}_{i}\in\mathbb{R}^{2k\times k}\text{ has orthonormal columns,}$
$\bm{V}$ $=\operatorname{blockdiag}\left(\bm{V}_{1},\ldots,\bm{V}_{2^{L}}\right)\quad\text{where }\bm{V}_{i}\in\mathbb{R}^{2k\times k}\text{ has orthonormal columns, and}$
@@ 3 Near-optimal greedy HSS approximation
3:Partition $\bm{A}^{(\ell+1)}=\{\bm{A}^{(\ell+1)}_{i,j}\}_{i,j}^{2^{\ell}}$ into blocks of size $\bm{A}_{i,j}^{(\ell+1)}\in\mathbb{R}^{2k\times 2k}$.
4:for $i=1,2,\ldots,2^{\ell}$ do
@@
$\hskip 7.39998pt\leq\Gamma_{\textup{d}}\cdot\Big{[}\|r_{i}(\bm{A}^{(\ell+1)})-\bm{U}_{i}^{(\ell)}(\bm{U}_{i}^{(\ell)})^{\mathsf{T}}r_{i}(\bm{A}^{(\ell+1)})\|_{\mathsf{F}}^{2}+\|c_{i}(\bm{A}^{(\ell+1)})-c_{i}(\bm{A}^{(\ell+1)})\bm{V}_{i}^{(\ell)}(\bm{V}_{i}^{(\ell)})^{\mathsf{T}}\|_{\mathsf{F}}^{2}\Big{]}.$
Then, [^3] We are not aware of any prior result guaranteeing the existence of an optimal HSS approximation. We therefore include a proof that an optimal HSS approximation always exists; see Appendix D.
(3) $\mathbb{E}\Bigl{[}\|\bm{A}-\bm{B}\|_{\mathsf{F}}^{2}\Big{]}\leq\left(\Gamma_{\textup{r}}+\Gamma_{\textup{c}}\right)(1+\Gamma_{\textup{d}})L\cdot{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\min\limits_{\bm{C}\in\operatorname{HSS}(L,k)}}\|\bm{A}-\bm{C}\|_{\mathsf{F}}^{2}.$
As discussed, Algorithm 1 can be implemented in many different ways [43, 51, 52, 75, 74]. We highlight some important examples below:
Explicit matrices: If $\bm{A}$ is dense and stored in memory, the simplest way to implement Algorithm 1 is to use the truncated SVD: explicitly form $\bm{A}^{(\ell+1)}$ and let $\bm{U}_{i}^{(\ell)}$ and $\bm{V}_{i}^{(\ell)}$ be the top $k$ left and right singular vectors of $r_{i}(\bm{A}^{(\ell+1)})$ and $c_{i}(\bm{A}^{(\ell+1)})$, respectively, and let $\bm{D}_{i,i}^{(\ell)}={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bm{A}_{i,i}^{(\ell+1)}}$. This implementation requires $O(N^{2}k)$ operations. [^4] Each iteration of the inner loop in lines 4-8 requires computing the SVD of two matrices of sizes $2k\times(2^{\ell+1}k-2k)$ and $(2^{\ell+1}k-2k)\times 2k$. Hence, at level $\ell$ the total cost of the inner loop is $O\left((2^{\ell+1})^{2}k^{3}\right)$. Computing $\bm{A}^{(\ell)}$ in line 9 also requires $O\left((2^{\ell+1})^{2}k^{3}\right)$ operations due to the block-diagonal structure of $\bm{U}^{(\ell)},\bm{V}^{(\ell)},$ and $\bm{D}^{(\ell)}$. Thus, each level $\ell$ requires $O\left((2^{\ell+1})^{2}k^{3}\right)$ operations. Summing over $\ell=L,\ldots,1$ and using $N=2^{L+1}k$ gives an overall computational cost of Algorithm 1 of $O(4^{L+1}k^{3})=O(N^{2}k)$ operations. Since at each level we compute the optimal subspaces of each block-row and block-column, in the context of Theorem 3.1, this yields $\Gamma_{\textup{r}}=\Gamma_{\textup{c}}=1$ and $\Gamma_{\textup{d}}=0$. Consequently, for this implementation of Algorithm 1, the output satisfies $\|\bm{A}-\bm{B}\|_{\mathsf{F}}^{2}\leq 2L{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\min\limits_{\bm{C}\in\operatorname{HSS}(L,k)}}\|\bm{A}-\bm{C}\|_{\mathsf{F}}^{2}.$
Implicit matrices: When $\bm{A}$ can only be accessed using matvecs, we leverage the “block-nullification” idea of [43] to implement Algorithm 1. This approach is described in Section 4, where we prove in Theorem 4.1 that the output of this algorithm satisfies $\mathbb{E}\|\bm{A}-\bm{B}\|_{\mathsf{F}}^{2}\leq O(L){\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\min\limits_{\bm{C}\in\operatorname{HSS}(L,k)}}\|\bm{A}-\bm{C}\|_{\mathsf{F}}^{2},$ yielding our main result.
@@
Remark 3.2.
One can also prove a deterministic version of Theorem 3.1. In that case, the assumptions would be deterministic bounds (rather than in expectation) on the quality of our approximations to the right and left subspaces of $r_{i}(\bm{A}^{(\ell+1)})$ and $c_{i}(\bm{A}^{(\ell+1)})$ at each level $\ell$, and the expectation in Eq. 3 would be removed. The statement and proof is otherwise identical. However, we emphasize that our main result is the stated probabilistic version for several reasons. First, hierarchical matrix algorithms generally work level-by-level, and Theorem 3.1 reflects that an approximation at one level can be correlated to those at all prior levels. This dependence between levels is one of the key difficulties in analyzing hierarchical matrix techniques. Secondly, we extend this result to the matvec access model. Matvec algorithms for hierarchical matrix approximation must be probabilistic in nature because they rely on sketching for low-rank approximation, which cannot be done deterministically.